pacman25:

If u = ❬6, –9❭ and v = ❬–24, 36❭ with an angle θ between the vectors, are u and v parallel or orthogonal? Explain. The vectors are parallel because cos θ = −1. The vectors are parallel because u • v = 0. The vectors are orthogonal because u • v = 0. The vectors are orthogonal because cos θ = −1.

pacman25:

@jhonyy9 could you help me out dawg?

Sherixn:

1. Use the dot product 2. Find the magnitude of both vectors The formula is arccos((v dot u)/(||v||*||u||)

pacman25:

so could you walk me through it @sherixn

Sherixn:

Sure! So first find u dot v which is defined as u * v = (u1)(v1) + (u2)(v2)

Sherixn:

which would be 6(-24) + -9(36)

Sherixn:

This equals -468 The magnitude formula is ||v|| = \[\sqrt{a^2+b^2}\] Same for ||u||= \[\sqrt{a^2+b^2}\]

Sherixn:

||v|| = sqrt[6^2+(-9)^2] ||u|| = sqrt[(-24)^2+36^2]

Sherixn:

||v||=sqrt(117) ||u||= 4sqrt(117)

Sherixn:

Now we take the inverse cosine of -468/sqrt(117)*4sqrt(117)

Sherixn:

This equals 180

Sherixn:

When the angle is 180, cos = -1 (you can look at the unit circle of reference)

Sherixn:

So the answer is A. "The vectors are parallel because cos θ = −1."

pacman25:

thank you, sorry for the late reply i was busy @sherixn

Sherixn:

No problem, I figured @pacman25 (: